OpenStudy (anonymous):

A line has slope 5/6 and y–intercept –3. Which is the equation of the line? A line has slope and y–intercept –3. Which is the equation of the line?

OpenStudy (anonymous):

y=mx+c

OpenStudy (igreen):

Put in the form of y = mx + b Where m = slope, and b = y-intercept. y = mx + b Plug in Slope(5/6) and the y-intercept(-3) y = 5/6x - 3

OpenStudy (anonymous):

where m is slope

OpenStudy (igreen):

@CrazyCountryGirl

OpenStudy (anonymous):

@iGreen I was partially right. ;) I thought it was y=5/6x+3 I see where I went wrong though. Thanks!

OpenStudy (igreen):

Yes, it could be y = 5/6x + -3, or y = 5/6x - 3, but not y = 5/6x + 3.